the perimeter of a standard-size regular rug is 36ft. the length is 2 feet longer than the width. find the dimension.
Answer by rfer(16322)

2x+2(x+2)=36
2x+2x+4=36
4x=36-4
4x=32
x=32/4
x=8ft
x+2=10ft